A car starts from rest and accelerated uniformly for 5s until it attains a velocity of 3m-1. It then travels with uniform velocity for 15s before decelerates uniformly to rest in 10s.

  1. Sketch a graph of motion
  2. Using the graph above , calculate the:
  • acceleration during the first 5s.
  • Deceleration during the last 10s and total distance covered through the motion
1
Expert's answer
2021-08-29T16:55:26-0400
"a_1=\frac{3}{5}=0.6\frac{m}{s^2}\\
a_2=\frac{3}{10}=0.3\frac{m}{s^2}\\S=0.5(3)(5)+(3)(15)+0.5(3)(10)=67.5\ m"